# DELTA: A DAG-aware Efficient OCS Logical Topology Optimization Framework for AIDCs

### 30 Mar 2026 v1

@@ I Introduction
Driven by the vision of Artificial General Intelligence (AGI) and the scaling laws of LLMs [24, 18], AIDCs are undergoing rapid expansion. However, the evolution of computational demand has outpaced the iteration of underlying network equipment, making communication an increasingly prominent bottleneck for cluster scaling [58, 10, 13, 4, 53]. Existing multi-tier electrical Clos networks that support massive GPU interconnections struggle to bridge this gap. Specifically, as endpoint GPU communication rates surge, relying exclusively on electrical packet switches (EPS) not only introduces severe bottlenecks in power consumption [42, 46, 4, 5, 2] and latency [5, 2, 8], but the frequent upgrades to match these endpoint rates also incur prohibitive infrastructure churn [41, 47, 33]. Consequently, the industry is introducing OCS—leveraging their low power, low latency, and data-rate transparency—to replace the EPS in original Core or Spine layers, driving the emergence of OCS-AIDCs that enable efficient inter-Pod optical connectivity while retaining flexible intra-Pod electrical switching [41, 47, 51, 33, 22, 61].
However, unlike the flexible any-to-any connectivity of EPS, the coarse-grained, point-to-point nature of OCS struggles to accommodate the dynamic communication patterns in LLM training. Furthermore, while an OCS can reconfigure optical circuits within tens of milliseconds [47, 33, 31], the subsequent initialization of associated components—such as transceivers, NICs, and EPSs (update the routing tables in order)—requires several seconds [56, 57, 31]. Given that mainstream LLM training iterations typically span only a few seconds [37, 20, 21, 12, 30], intra-iteration topology reconfiguration is impractical. Under this constraint, maintaining a static topology throughout an iteration is a pragmatic strategy for current deployments [33, 51, 22, 61]. Consequently, a priori engineering of a logical topology (i.e., the number of optical circuits allocated between Pod pairs [59, 41, 14]) tailored to absorb time-varying LLM traffic is critical for mitigating OCS-induced communication bottlenecks.
@@ II-B Challenges of Optimizing Logical Topology for OCS-AIDC
The transition from the any-to-any connectivity of EPS (Fig. 2(a)) to the rigid point-to-point nature of OCS creates bottlenecks when destination-divergent flows (e.g., orthogonal PP and DP traffic) compete for limited optical exit ports (Fig. 2(b)). Under such hardware constraints, topology construction is essentially a constrained resource allocation problem. While utilizing the aforementioned opportunities for this problem is promising, capturing these benefits is non-trivial due to two distinct challenges in encoding the dynamic LLM training traffic patterns into the optimization formulation.
(a) Any-to-any inter-pod paths established by the EPS.
(b) Point-to-point inter-pod paths established by the OCS.
Figure 2: Comparison of inter-pod connection patterns established by OCS and EPS for the multi-pod LLM training setup in Fig. 1.
@@
C2: Challenge in encoding diverse and decision-dependent communication demands into topology optimization. Encoding the time-varying bandwidth requirements of LLM workloads (F2 and F3) into the optimization formulation presents two algorithmic hurdles. First, communication demands driven by different parallelization strategies exhibit distinct temporal behaviors, such as the stepped bandwidth demand envelopes of DP traffic versus the uniform, rectangular demand of PP traffic. Constructing a topology solely based on a traffic matrix fails to capture these temporal dynamics, thereby missing the temporal multiplexing opportunities created by the pipeline-induced staggering of DP flows (i.e., failing to exploit O3 and O4). Second, a circular dependency exists between topology decisions and traffic profiles. Specifically, any adjustment to the logical topology inherently alters transmission rates, which in turn stretches or compresses the actual flow durations. This non-linear coupling invalidates static optimization approaches, requiring a solver capable of dynamically tracking how topological decisions reshape the temporal demand landscape.
@@ II-C Related Work
Introducing OCS into AI data centers. Recent studies have begun integrating OCS into AI data centers to optimize AI workload training time while reducing network costs [35, 48, 9, 55, 34, 31, 43, 1, 25, 51]. Early works [35, 48] optimized OCS reconfiguration and multiplexed the optical circuit by exploiting the interleaved compute-communicate phases of distributed training, but focused exclusively on DP parameter synchronization. To further accelerate this synchronization process, [9, 55, 34] jointly optimize the schedule of optical circuit and in-network computing resources. However, such DP-centric and hardware-dependent designs lack the generality required for modern LLM training. Other studies [31, 43, 1] explore architectures where GPUs are directly connected to OCS, reconfiguring the topology based on traffic pattern predictions to optimize the All-to-All or All-Reduce communications required for EP or TP/DP. Yet, this approach is incompatible with the prevailing architectural paradigm of scaling clusters by interconnecting multiple pods via OCS. While the methods proposed in [25, 51] support OCS-based pod-level interconnection and are not restricted to specific workloads, their decision-making processes still rely on aggregated traffic matrices and fail to incorporate the iteration time of LLM training as an optimization objective. Consequently, they do not resolve the previously mentioned challenges C1 and C2, which may lead to sub-optimal topological decisions—a critical limitation that our work specifically aims to overcome.
@@ III-A Encoding LLM Training Traffic Dynamics via the Computation-Communication DAG
where $\mathcal{E}_{\text{crit}}^{\text{inter-pod}}$ denotes the inter-pod tasks on $\mathcal{E}_{\text{crit}}$. Eq. (2) provides a reduced-complexity model for LLM training iteration time, which will be incorporated into the mathematical formulation of the topology optimization problem in the subsequent section.
@@ III-B MILP Formulation with Variable-Length Time Interval for Logical Topology Optimization
To overcome this computational bottleneck, we pioneer a variable-length time interval modeling approach inspired by the methodology of Discrete Event Simulation (DES), as illustrated in Fig. 4. In a DES paradigm, the system state—specifically, the task activation state ($y_{m,k}$) and flow rates (the ratio of data volume $w_{m,k}$ to interval duration $\Delta_{k}$)—remains entirely static between state-transition events, which in our context correspond to the initiation ($S_{m}$) and completion ($C_{m}$) of communication tasks. Therefore, instead of uniformly slicing the time horizon, we can derive the temporal discretization points ($t_{k}$) solely from the occurrences of these events, dynamically partitioning the timeline into a sequence of $K$ variable-length intervals. [^1] Theoretically, because each task contributes exactly two state-transition timestamps (initiation and completion), setting $K\leq 2|\mathcal{M}|-1$ (where $\mathcal{M}$ is the set of all tasks) guarantees capturing all state transitions. In practice, $K$ is profiled from a baseline simulation and is typically much smaller, as synchronized tasks (e.g., equivalent communication tasks executing concurrently within separate DP replicas in identical network environments) share identical event timestamps.
@@
  • $\mathcal{D}$: The set of temporal dependencies among communication tasks. The element $(m_{\text{pre}},m,\delta_{m_{\text{pre}}\to m})\in\mathcal{D}$ indicates that the successor task $m$ can only start after a rigid time interval of $\delta_{m_{\text{pre}}\to m}$ following the completion of the predecessor task $m_{\text{pre}}$. Note that incorporating the a priori knowledge of $\mathcal{M}$ and $\mathcal{D}$ exploits O1 for proactive topology engineering and O3 for DP flow-staggering-aware topology optimization.
@@
IV Acceleration of MILP Solving and Design of Heuristic Algorithms
Despite the substantial simplification achieved through the reduced inter-pod communication DAG and variable-length time interval modeling, directly solving the proposed MILP remains computationally intractable for LLM training workloads spanning multiple Pods and thousands of GPUs. To overcome this, this section introduces a dual-track acceleration approach. First, we incorporate domain-specific strategies to reduce the MILP scale. In parallel, we develop a fast heuristic search algorithm for efficient hot-starting. Collectively, these strategies guarantee rapid convergence to a high-quality solution within minutes, even at the thousand-GPU scale.
IV-A Strategies for Accelerating MILP Solving
@@
(a) Active intervals of tasks under a baseline network in practice.
@@
While the preceding topological sorting establishes boundaries based on whole-graph-wide longest paths, relying solely on it yields overly conservative search spaces (detailed in Appendix B). From a local-task-centric perspective, as observed from Fig. 3(b) and Fig. 5(a), the dependencies in $\mathcal{D}$ strictly constrain the execution of intermediate tasks: a task $m$ cannot overlap with the state transitions of its predecessors or successors. Driven by this strict causal ordering, an intermediate task typically monopolizes a complete time interval, leaving minimal room for shifts within the discrete interval sequence and rendering its active index range highly deterministic. Consequently, a lightweight baseline simulation (Fig. 5(a)) can approximate this range. Exploiting this predictability, we extract index bounds $\tilde{k}^{\text{start}}_{m},\tilde{k}^{\text{end}}_{m}]$ for each intermediate task $m$ from the profiling results (to accommodate potential index encroachments among tasks sharing identical topological orders, we introduce a slight redundancy margin when calculating $\tilde{k}^{\text{start}}_{m},\tilde{k}^{\text{end}}_{m}]$). By integrating these anchoring bounds with longest-path-derived boundaries, we tightly confine the allowable active indices, further pruning the mkm-k search space as illustrated by the dark green regions in Fig. 5(b). The complete algorithm workflow is presented in Alg. 1. Since most tasks are intermediate tasks (Fig. 3(b)), bounding their indices significantly tightens the entire MILP scale. Through a cascade effect, this tightening further prunes the indices of DP tasks. Ultimately, the combined restriction on both intermediate and DP tasks reduces the scale of task-time decision variables from $\mathcal{O}(|\mathcal{M}|^{2})$ to $\mathcal{O}(|\mathcal{M}|)$.
@@
IV-B Simulation-Based Heuristic Search Algorithm
Although the accelerated MILP in Section IV-A efficiently achieves joint topology and flow-rate optimization, broader scenarios—like GPU-resource allocation co-optimization in multi-tenant clusters—demand frequent invocations of the topology optimizer. To this end, we propose a fast, simulation-based heuristic search algorithm (DELTA-Fast). The key idea is to decouple the search space by offloading the constraint-solving process of the cumbersome task-time dimensions (i.e., $y_{m,k}$, $w_{m,k}$, and $\Delta_{k}$) to a lightweight DES engine, thereby avoiding computationally expensive joint searches across all variables. Specifically, an outer-loop genetic algorithm optimizes the logical topology ($x_{i,j}$), while an inner-loop DES chronologically simulates tasks based on DAG dependencies. Through a single simulation pass, it implicitly determines valid values for these temporal and task-state variables, naturally satisfying all constraints. Although this decoupling sacrifices the dynamic joint optimization of flow rates (degenerating to a fair-sharing mechanism), it effectively mitigates computational complexity, enabling the rapid generation of high-quality topology candidates for upper-level co-optimization frameworks. Furthermore, since the DES execution trace is isomorphic to the MILP’s event-driven formulation, the simulated results seamlessly map to an initial feasible solution, providing an efficient hot-start for the MILP. To further accelerate the search process, we design DELTA-Fast around two core components: a DAG-aware search space pruning strategy and a domain-adapted genetic algorithm.
@@ IV-B Simulation-Based Heuristic Search Algorithm
2) Simulation-Based Domain-Adapted Genetic Algorithm:

### 02 Apr 2026 v2

# Beyond Traffic Matrix: DELTA – A DAG-Aware OCS Logical Topology Optimization for AIDCs

@@ I Introduction
Driven by the vision of Artificial General Intelligence (AGI) and the scaling laws of LLMs [24, 18], AIDCs are undergoing rapid expansion. However, the evolution of computational demand has outpaced the iteration of underlying network equipment, making communication an increasingly prominent bottleneck for cluster scaling [60, 10, 13, 4, 55]. Existing multi-tier electrical Clos networks that support massive GPU interconnections struggle to bridge this gap. Specifically, as endpoint GPU communication rates surge, relying exclusively on electrical packet switches (EPS) not only introduces severe bottlenecks in power consumption [43, 48, 4, 5, 2] and latency [5, 2, 8], but the frequent upgrades to match these endpoint rates also incur prohibitive infrastructure churn [42, 49, 34]. Consequently, the industry is introducing OCS—leveraging their low power, low latency, and data-rate transparency—to replace the EPS in original Core or Spine layers, driving the emergence of OCS-AIDCs that enable efficient inter-Pod optical connectivity while retaining flexible intra-Pod electrical switching [42, 49, 44, 53, 34, 22, 63].
However, unlike the flexible any-to-any connection of EPS, the coarse-grained, point-to-point connectivity of OCS struggles to accommodate the dynamic traffic patterns in LLM training. Furthermore, while an OCS can reconfigure its optical circuits within tens of milliseconds [49, 34, 32], the subsequent initialization of transceivers, NICs, and EPSs (updating the routing tables in order)—requires several seconds [58, 59, 32]. Given that mainstream LLM training iterations typically span only a few seconds [38, 20, 21, 12, 31], intra-iteration topology reconfiguration is impractical. Under this constraint, maintaining a static topology throughout an iteration is a pragmatic strategy for current deployments [34, 53, 22, 63]. Consequently, a priori engineering of a logical topology (i.e., the number of optical circuits allocated between Pod pairs [61, 42, 14]) tailored to absorb time-varying LLM traffic is critical for mitigating OCS-induced communication bottlenecks.
@@ II-B Challenges of Optimizing Logical Topology for OCS-AIDC
The transition from the any-to-any connections of EPS (Fig. 2(a)) to the rigid point-to-point connectivity of OCS introduces bottlenecks, as flows with diverging destinations (e.g., orthogonal PP and DP traffic) contend for limited optical exit ports (Fig. 2(b)). Under such hardware constraints, topology construction is essentially a constrained resource allocation problem. While utilizing the aforementioned opportunities for this problem is promising, capturing these benefits is non-trivial due to two distinct challenges in encoding the dynamic LLM training traffic patterns into the optimization formulation.
(a) Any-to-any inter-pod paths established by the EPSs.
(b) Point-to-point inter-pod paths established by the OCSs.
Figure 2: Comparison of EPS and OCS inter-pod connection patterns for accommodating the LLM training communication demands in Fig. 1 For visual clarity, the multiple EPSs/OCSs within the core layer are depicted as a single switching plane.
@@
C2: Challenge in encoding diverse communication demands and formulating an optimization algorithm to resolve their non-linear coupling with topological decisions. Encoding the time-varying bandwidth requirements of LLM workloads (F2 and F3) into the optimization formulation presents two algorithmic hurdles. First, communication demands driven by different parallelization strategies exhibit distinct temporal behaviors, such as the stepped bandwidth demand envelopes of DP traffic versus the uniform, rectangular demand of PP traffic. Constructing a topology solely based on a traffic matrix fails to capture these temporal dynamics, thereby missing the temporal multiplexing opportunities created by the pipeline-induced staggering of DP flows (i.e., failing to exploit O3 and O4). Second, a circular dependency exists between topology decisions and traffic profiles. Specifically, any adjustment to the logical topology inherently alters transmission rates, which in turn stretches or compresses the actual flow durations. This non-linear coupling renders open-loop static optimization ineffective, necessitating a solver capable of dynamically capturing how topological configurations reshape temporal traffic profiles.
@@ II-C Related Work
Introducing OCS into AI data centers. Recent studies have begun integrating OCS into AI data centers to optimize AI workload training time while reducing network costs [36, 50, 9, 57, 35, 32, 45, 1, 25, 53]. Early works [36, 50] optimized OCS reconfiguration and multiplexed the optical circuit by exploiting the interleaved compute-communicate phases of distributed training, but focused exclusively on DP parameter synchronization. To further accelerate this synchronization process, [9, 57, 35] jointly optimize the schedule of optical circuits and in-network computing resources. However, such DP-centric and hardware-dependent designs lack the generality required for modern LLM training. Other studies [32, 45, 1] explore architectures where GPUs are directly connected to OCS, reconfiguring the topology based on traffic pattern predictions to optimize the All-to-All or All-Reduce communications required for EP or TP/DP. Yet, this approach is incompatible with the prevailing architectural paradigm of scaling clusters by interconnecting multiple pods via OCS. While the methods proposed in [25, 53, 30] support OCS-based pod-level interconnection and are not restricted to specific workloads, their decision-making processes still rely on aggregated traffic matrices and fail to incorporate the iteration time of LLM training as an optimization objective. Consequently, they do not resolve the previously mentioned challenges C1 and C2, which may lead to sub-optimal topological decisions—a critical limitation that our work specifically aims to overcome.
@@ III-A Encoding LLM Training Traffic Dynamics via the Computation-Communication DAG
where $\mathcal{E}_{\text{crit}}^{\text{inter-pod}}$ denotes the inter-pod tasks on $\mathcal{E}_{\text{crit}}$. Eq. (2) provides a reduced-complexity model for LLM training iteration time, which will be implicitly incorporated into the formulation of the MILP in the subsequent section.
@@ III-B MILP Formulation with Variable-Length Time Interval for Logical Topology Optimization
To overcome this computational bottleneck, we pioneer a variable-length time interval modeling approach inspired by the methodology of Discrete Event Simulation (DES), as illustrated in Fig. 4. In a DES paradigm, the system state—specifically, the task activation state ($y_{m,k}$) and flow rates (the ratio of data volume $w_{m,k}$ to interval duration $\Delta_{k}$)—remains entirely static between state-transition events, which in our context correspond to the initiation ($S_{m}$) and completion ($C_{m}$) of communication tasks. Therefore, instead of uniformly slicing the time horizon, we formulate the time interval durations ($\Delta_{k}$) as decision variables and derive the temporal discretization points ($t_{k}$) from the occurrences of these events, dynamically partitioning the timeline into a sequence of $K$ variable-length intervals. [^1] Theoretically, because each task contributes exactly two state-transition timestamps (initiation and completion), setting $K\leq 2|\mathcal{M}|-1$ (where $\mathcal{M}$ is the set of all tasks) guarantees capturing all state transitions. In practice, $K$ is profiled from a baseline simulation and is typically much smaller, as synchronized tasks (e.g., equivalent communication tasks executing concurrently within separate DP replicas in identical network environments) share identical event timestamps.
@@
  • $\mathcal{D}$: The set of temporal dependencies among communication tasks. The element $(m_{\text{pre}},m,\delta_{m_{\text{pre}}\to m})\in\mathcal{D}$ indicates that the successor task $m$ can only start after a rigid time interval of $\delta_{m_{\text{pre}}\to m}$ following the completion of the predecessor task $m_{\text{pre}}$. Note that incorporating the a priori knowledge of $\mathcal{M}$ and $\mathcal{D}$ (the DAG of LLM training) exploits O1 for proactive topology engineering and O3 for DP flow-staggering-aware topology optimization.
@@
IV Dual-Track Acceleration by MILP Search Space Pruning and DES-Accelerated Heuristics
Despite the substantial simplification achieved through the reduced inter-pod communication DAG and variable-length time interval modeling, directly solving the proposed MILP remains computationally intractable for LLM training workloads spanning multiple Pods and thousands of GPUs. To overcome this, this section introduces a dual-track acceleration approach. First, we incorporate domain-specific strategies to reduce the MILP scale. In parallel, we develop a DES-accelerated heuristic search algorithm for fast topology search and hot-starting. Collectively, these strategies guarantee rapid convergence to a high-quality solution within minutes, even at the thousand-GPU scale.
IV-A Strategies for Pruning MILP Search Space
@@
(a) Active intervals of tasks under a baseline simulation.
@@
While the preceding topological sorting establishes boundaries based on whole-graph-wide longest paths, relying solely on it yields overly conservative search spaces (detailed in Appendix B). From a local, task-centric perspective, as observed from Fig. 3(b) and Fig. 5(a), the dependencies in $\mathcal{D}$ strictly constrain the execution of intermediate tasks: a task $m$ cannot overlap with the state transitions of its predecessors or successors. Driven by this strict causal ordering, an intermediate task typically monopolizes a whole time interval, leaving minimal room for shifts within the discrete interval sequence and rendering its active index range highly deterministic. Consequently, a lightweight baseline simulation (Fig. 5(a)) can approximate this range. Exploiting this predictability, we extract index bounds $\tilde{k}^{\text{start}}_{m},\tilde{k}^{\text{end}}_{m}]$ from the profiling results to act as reference anchors that localize the execution of each intermediate task $m$. By integrating these anchoring bounds with the longest-path-derived boundaries, we tightly confine the allowable active indices, further pruning the mkm-k search space as illustrated by the dark green regions in Fig. 5(b). The complete algorithm workflow is presented in Alg. 1. Since most tasks are intermediate tasks (Fig. 3(b)), bounding their indices significantly tightens the entire MILP scale. Through a cascade effect, this tightening further prunes the indices of DP tasks. Ultimately, the combined restriction on both intermediate and DP tasks reduces the scale of task-time decision variables from $\mathcal{O}(|\mathcal{M}|^{2})$ to $\mathcal{O}(|\mathcal{M}|)$.
@@
IV-B DES-Accelerated Heuristics for Fast Topology Search
Although the accelerated MILP in Section IV-A efficiently achieves joint topology and flow-rate optimization, broader scenarios—like GPU-resource allocation co-optimization in multi-tenant clusters—demand frequent invocations of the topology optimizer. To this end, we propose a fast, DES-accelerated heuristic search algorithm (DELTA-Fast). The key idea is to decouple the search space by offloading the constraint-solving process of the cumbersome task-time dimensions (i.e., $y_{m,k}$, $w_{m,k}$, and $\Delta_{k}$) to a lightweight DES engine, thereby avoiding computationally expensive joint searches across all variables. Specifically, an outer-loop genetic algorithm optimizes the logical topology ($x_{i,j}$), while an inner-loop DES chronologically simulates tasks based on DAG dependencies. Through a single simulation pass, it determines valid values for these temporal and task-state variables, and implicitly satisfies all constraints. Although this decoupling sacrifices the dynamic joint optimization of flow rates (degenerating to a fair-sharing mechanism), it effectively mitigates computational complexity, enabling the rapid generation of high-quality topology candidates for upper-level co-optimization frameworks. Furthermore, since the DES execution trace is isomorphic to the MILP’s event-driven formulation, the simulated results seamlessly map to an initial feasible solution, providing an efficient hot-start for the MILP. To further accelerate the search process, we design DELTA-Fast around two core components: a DAG-aware search space pruning strategy and a domain-adapted genetic algorithm.
@@ IV-B DES-Accelerated Heuristics for Fast Topology Search
2) DES-Accelerated Domain-Adapted Genetic Algorithm: